The number of potholes in any given 1-mile stretch of freeway pavement in Pennsylvania has a Normal distribution. This distribution has a mean of 60 and a standard deviation of 6. Using the Empirical Rule, what is the approximate percentage of 1-mile long roadways with potholes numbering between 54 and 78? do not round your answer.
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 60
std deviation   =σ= 6.000
probability =P(54<X<78)=P((54-60)/6)<Z<(78-60)/6)=P(-1<Z<3)=99.85-16.0 =83.85 %